# GUI Agents: A Survey

### v1

@@ 1 Introduction
Large Foundation Models (LFMs) have significantly transformed both the landscape of AI research and day-to-day life (Bommasani et al., 2022; Kapoor et al., 2024; Schneider et al., 2024; Naveed et al., 2023; Wang et al., 2024d). Recently, we have witnessed a paradigm shift from using LFMs purely as conversational chatbots (Touvron et al., 2023; Chiang et al., 2023; Dam et al., 2024) to employing them for performing actions and automating useful tasks (Wang et al., 2024b; Zhao et al., 2023; Yao et al., 2023; Shinn et al., 2023; Shen et al., 2024b; Cheng et al., 2024c). In this direction, one approach stands out: leveraging LFMs to interact with digital systems, such as desktops and mobile phones, or software applications such as a web browser, through Graphical User Interfaces (GUIs) in the same way humans do—for example, by controlling the mouse and keyboard to interact with visual elements displayed on a device’s monitor Iong et al. (2024); Hong et al. (2023); Lu et al. (2024); Shen et al. (2024a).
This approach holds great potential, as GUIs are ubiquitous across almost all computer devices that humans interact with in their work and daily lives. However, deploying LFMs in such environments poses unique challenges, such as dynamic layouts, diverse graphical designs across different platforms, and grounding issues—for instance, fine-grained recognition of elements within a page that are often small, numerous, and scattered Liu et al. (2024b). Despite these challenges, many early efforts have shown significant promise (Lin et al., 2024; Cheng et al., 2024a), and growing interest from major players in the field is becoming evident [^1] Anthropic, Google DeepMind, OpenAI.
@@
Organization of this Survey. We begin our survey by clearly defining the term “GUI Agent,” followed by a traditional RL formalism of GUI Agent tasks in Section 2. We then summarize different datasets and environments in Section 3 to provide readers a clearer picture of the kinds of problem settings currently available. We summarize various GUI Agent architectural designs in Section 4, followed by different ways of training them in Section 5. Lastly, we discuss open problems and future prospects of GUI Agent research in Section 6.
@@ 2 Preliminaries
GUI Agent involves an agent interacting with an environment in a sequential manner. The environment can generally be modeled as a Partially Observable Markov Decision Process (POMDP), defined by a tuple $(\mathcal{U},\mathcal{A},\mathcal{S},\mathcal{O},T)$, where $\mathcal{U}$ is the task space, $\mathcal{A}$ is the action space, $\mathcal{S}$ is the state space, $\mathcal{O}$ is the observation space, and $T:\mathcal{S}\times\mathcal{A}\to\mathcal{P}(\mathcal{S})$ is a state transition function mapping a state-action pair to a probability distribution over subsequent states.
A GUI Agent is a mapping $\pi:\mathcal{O}\to\mathcal{A}$. Given a task $u\in\mathcal{U}$, the agent proceeds through a sequence of actions to complete the task. At each step $t$, based on the available observation $o\in\mathcal{O}$, the agent $\pi$ must predict the next action $a\in\mathcal{A}$. The environment then transitions to a new state $s^{\prime}\in\mathcal{S}$ according to $T$. Depending on the design of the environment, the agent may receive a reward $r=R(s,a,s^{\prime})$, where $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ is a reward function.
@@ 3 Benchmarks
When evaluating GUI Agents, it is crucial to distinguish between an environment and a dataset. A dataset is a static collection of data point, where each consists of several input features (e.g., a question, a screenshot of the environment, or the current state of the environment) and some output features (e.g., correct answers or actions to be taken). A dataset remains unchanged throughout the evaluation process. In contrast, an environment is an interactive simulation that represents a real-world scenario of interest. A GUI environment includes the GUI interface of a mobile phone or a desktop. Unlike datasets, environments are dynamic—actions taken within the environment can alter its state, hence, allowing modeling the problem as Markov Decision Processes (MDPs) or Partially Observable MDPs (POMDPs), with defined action, state, and observation spaces, and a state transition function.
@@ 3.1 Static Datasets
3.1.1 Closed-World Datasets
@@
Evaluation on static open-world datasets balances the ease of setting up an evaluation setting with realism since the agents interact with real-world websites. However, due to the nature of real-world websites, they are often unpredictable and prone to changes, which makes it more challenging to reproduce and compare with prior methods.
@@ 3.2 Interactive Environments
Closed-world interactive environments provide controlled and reproducible settings for evaluating agent capabilities. MiniWoB offers synthetic web tasks requiring interactions with webpages using mouse and keyboard inputs (Shi et al., 2017). It focuses on fundamental skills like button clicking and form filling, providing a baseline for evaluating low-level interaction. CompWoB extends MiniWoB with compositional tasks, requiring agents to handle multi-step workflows and generalize across task sequences (Furuta et al., 2023). This introduces dynamic dependencies that reflect real-world complexity. WebShop simulates e-shopping tasks that challenge agents to navigate websites, process instructions, and make strategic decisions (Yao et al., 2022). WebArena advances realism with self-hosted environments across domains like e-commerce and collaborative tools, requiring agents to manage long-horizon tasks (Zhou et al., 2023b). VisualWebArena adds multimodal challenges, integrating visual and textual inputs for tasks like navigation and object recognition (Koh et al., 2024a). Shifting to enterprise settings, WorkArena evaluates agent performance in complex UI environments, focusing on knowledge work tasks in ServiceNow platform (Drouin et al., 2024). ST-WebAgentBench incorporates safety and trustworthiness metrics, assessing policy adherence and minimizing risky actions, critical for business deployment (Levy et al., 2024). Lastly, VideoWebArena introduces long-context video-based tasks, requiring agents to understand instructional videos and integrate them with textual and visual data to complete tasks. It emphasizes memory retention and multimodal reasoning (Jang et al., 2024).
@@ 3.3 Evaluation Metrics
Task Completion Metrics.
@@
The majority of benchmarks use task completion rate as the primary metric to measure GUI Agents’ performance. However, different papers define task completion differently. Success can be defined as whether an agent successfully stops at a goal state (Chen et al., 2024; Zhou et al., 2023b), with Zhou et al. (2023b) programmatically checking if the intended outcome has been achieved (e.g., a comment has been posted, or a form has been completed), or whether the returned results exactly match the ground truth labels (Shi et al., 2017; Yao et al., 2022; Koh et al., 2024a; Drouin et al., 2024; Levy et al., 2024; Mialon et al., 2023). Another approach is to measure success based on whether an agent completes all required subtasks (Lai et al., 2024; Zhang et al., 2024e; Pan et al., 2024; Furuta et al., 2023; Jang et al., 2024; Cheng et al., 2024b). This approach can be further extended to measure partial success, as shown in Zhang et al. (2024e). WebVoyager uses GPT-4V to automatically determine success based on the agent’s trajectory, reporting a high agreement rate of 85.3% with human judgments (He et al., 2024b). Instead of using a single final-state success metric, WebLINX measures an overall success rate based on aggregated turn-level success rates across tasks (Lù et al., 2024). The turn-level success rates are computed depending on the type of actions, e.g., Intersection Over Union (IoU) for click or submit actions, and F1 for say or textinput actions. Lastly, there are task-specific metrics to measure success, e.g., using ROUGE-L, F1 for open-ended generation (Liu et al., 2024b; Xu et al., 2024; Wornow et al., 2024), accuracy for multiple choice question tasks (Liu et al., 2024b), Precision and Recall for Standard Operating Procedure (SOP) validation (Wornow et al., 2024), and so on.
Intermediate Step Metrics.
@@
While the task completion rate is a straightforward single-numeric metric that simplifies comparing the overall performance of agents, it fails to provide clear insights into their specific behaviors. Although some fine-grained metrics measure step-wise performance, their scope remains limited. WebCanvas evaluates step scores using three distinct targets: URL Matching, which verifies whether the agent navigated to the correct webpage; Element Path Matching, which checks if the agent interacted with the appropriate UI element, such as a button or text box; and Element Value Matching, which ensures the agent inputted or extracted the correct values, such as filling a form or reading text. WebLINX uses an intent match metric to assess whether the predicted action’s intent aligns with the reference intent. Similarly, Mind2Web and MT-Mind2Web evaluate Element Accuracy by measuring the rate at which the agent selects the correct elements. These systems also measure the precision, recall, and F1 score for token-level operations, such as clicking or typing, and calculate the Step Success Rate, which reflects the proportion of individual task steps completed correctly. While step-wise evaluations provide more fine-grained insight into the agent’s performance, it is often challenging to collect reference labels at the step level while also providing enough flexibility to consider different paths to achieve the original tasks.
Efficiency, Generalization, Safety and Robustness Metrics.
@@ 4 GUI Agent Architectures
This section focuses on various architectural designs of a GUI Agent agent, which we categorize into four main types: (1) Perception: designs that enable the GUI Agent agent to perceive and interpret observations from its environment; (2) Reasoning: designs related to the cognitive processes of a GUI Agent agent, such as using an external knowledge base for long-term memory access or a world model of the environment to support other modules like planning; (3) Planning: designs related to decomposing a task into subtasks and creating a plan for their execution; and (4) Acting: mechanisms that allow the GUI Agent agent to interact with the environment, including representing actions in natural language using specific templates, JSON, or programming languages as action representations.
@@ 4.1 Perception
Unlike API-based agents that process structured, program-readable data, GUI agents must perceive and understand the on-screen environment that is designed for human consumption. This requires carefully chosen interfaces that allow agents to discover the location, identity, and properties of the interactive elements. Broadly, these perception interfaces can be categorized into four types: accessibility-based, HTML/DOM-based, screen-visual-based, and hybrid ones, with each offering different capabilities and posing distinct privacy and implementation considerations.
@@ 4.1.3 Screen-visual-based Interfaces
With advances in computer vision and multimodal LLM, agents can utilize screen-visual information, like screenshots, to perceive on-screen environment. OmniParser (Lu et al., 2024) utilizes an existing multimodal LLM (e.g., GPT-4V) to parse a screenshot into a structured representation of the UI elements. However, screen-visual-based perception introduces privacy concerns since entire screenshots may contain sensitive information. Additionally, computational overhead increases as models must handle high-dimensional image inputs. Despite these challenges, such interfaces are crucial for agents operating in environments where high-quality accessibility interfaces and DOM information are unavailable, or environments where dynamic or visual information is crucial, like image or video editing software.
@@ 4.1.4 Hybrid Interfaces
To achieve robust and flexible performance across diverse environments, many GUI agents employ a hybrid approach. These systems combine accessibility APIs, DOM data, and screen-visual information to form a more comprehensive understanding of the interface. Leading methods in GUI agent tasks, such as OS-Atlas(Wu et al., 2024b) and UGround (Gou et al., 2024), demonstrates that hybrid interfaces that combine visual and textual inputs can enhance performance. Hybrid interfaces based approaches also facilitate error recovery—when accessibility or DOM data are incomplete or misleading, the agent can fall back on screen parsing, and vice versa.
@@ 4.3 Planning
Planning involves decomposing a global task into multiple subtasks that progressively approach the goal state starting from an initial state Huang et al. (2024). Traditional planning methods, such as symbolic approaches and reinforcement learning, have significant limitations: symbolic methods require extensive human expertise to define rigid system rules and lack error tolerance Belta et al. (2007); Pallagani et al. (2022), while reinforcement learning demands impractical volumes of training data, often derived from costly environmental interactions Acharya et al. (2023). Recent advancements in LLM-powered agents offer a transformative alternative by positioning LLM-powered agents as the cognitive core for planning agents Huang et al. (2024). When equipping agents with GUIs as the medium, LLM-powered agents can directly interact with nearly all application domains and resources to enhance planning strategies. Based on what application domains/resources agents use for planning, we divide existing works into planning with internal and external knowledge.
@@ 4.3.1 Planning with Internal Knowledge
Planning with internal knowledge of GUI agents is to leverage the inherent knowledge to reason and think about the potential plans to fulfill the global task goals Schraagen et al. (2000). WebDreamer (Gu et al., 2024) uses LLMs to simulate the outcomes of the actions of each agent and then evaluate the result to determine the optimal plan at each step. MobA (Zhu et al., 2024) devises a two-level architecture to power the mobile phone management, with a high level for understanding user commands, tracking history memories and planning tasks, and a low level to act the planned module. Agent S (Agashe et al., 2024) introduces an experience-augmented hierarchical planning to perform complex computer tasks.

### v2

@@ 1 Introduction
Large Foundation Models (LFMs) have significantly transformed both the landscape of AI research and day-to-day life (Bommasani et al., 2022; Kapoor et al., 2024; Schneider et al., 2024; Naveed et al., 2023; Wang et al., 2024d). Recently, we have witnessed a paradigm shift from using LFMs purely as conversational chatbots (Touvron et al., 2023; Chiang et al., 2023; Dam et al., 2024) to employing them for performing actions and automating useful tasks (Wang et al., 2024b; Zhao et al., 2023; Yao et al., 2023; Shinn et al., 2023; Shen et al., 2024b; Cheng et al., 2024c). In this direction, one approach stands out: leveraging LFMs to interact with digital systems, such as desktops and mobile phones, or software applications such as a web browser, through Graphical User Interfaces (GUIs) in the same way humans do, for example, by controlling the mouse and keyboard to interact with visual elements displayed on a device’s monitor Iong et al. (2024); Hong et al. (2023); Lu et al. (2024); Shen et al. (2024a).
This approach holds great potential, as GUIs are ubiquitous across almost all computer devices that humans interact with in their work and daily lives. However, deploying LFMs in such environments poses unique challenges, such as dynamic layouts, diverse graphical designs across different platforms, and grounding issues, for instance, fine-grained recognition of elements within a page that are often small, numerous, and scattered Liu et al. (2024b). Despite these challenges, many early efforts have shown significant promise (Lin et al., 2024; Cheng et al., 2024a), and growing interest from major players in the field is becoming evident [^1] Anthropic, Google DeepMind, OpenAI.
@@
Organization of this Survey. We begin our survey by clearly defining the term “GUI Agent,” followed by a formal definition of GUI agent tasks in Section 2. We then summarize different datasets and environments in Section 3 to provide readers a clearer picture of the kinds of problem settings currently available. We summarize various GUI agent architectural designs in Section 4, followed by different ways of training them in Section 5. Lastly, we discuss open problems and future prospects of GUI agent research in Section 6.
@@ 2 Preliminaries
This section formally defines the term “GUI agent” and presents a formalization of GUI agent tasks.
@@
GUI agent tasks involve an agent interacting with an environment in a sequential manner. The environment can generally be modeled as a Partially Observable Markov Decision Process (POMDP) (Sondik, 1971; Hauskrecht, 2000), defined by a tuple $(\mathcal{U},\mathcal{A},\mathcal{S},\mathcal{O},T)$, where $\mathcal{U}$ is the task space, $\mathcal{A}$ is the action space, $\mathcal{S}$ is the state space (not fully observable to the agent), $\mathcal{O}$ is the observation space, and $T:\mathcal{S}\times\mathcal{A}\to\mathcal{P}(\mathcal{S})$ is a state transition function that maps a state-action pair to a probability distribution over subsequent states. A GUI agent is a policy $\pi:\Delta^{\mathcal{S}}\to\mathcal{A}$, where $\Delta^{\mathcal{S}}$ denotes the probability simplex over the state states. Most commonly, this is implemented using the entire history of past actions and observations. Given a task $u\in\mathcal{U}$, the agent proceeds through a sequence of actions to complete the task. At each step $t$, based on the history of past actions and observations, the policy $\pi$ selects the next action $a\in\mathcal{A}$. The environment then transitions to a new state $s^{\prime}\in\mathcal{S}$ according to $T$. Depending on the environment’s design, the agent may receive a reward $r=R(s,a,s^{\prime})$, where $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ is a reward function.
@@ 3 Benchmarks
When evaluating GUI agents, it is crucial to distinguish between an environment and a dataset. A dataset is a static collection of data point, where each consists of several input features (e.g., a question, a screenshot of the environment, or the current state of the environment) and some output features (e.g., correct answers or actions to be taken). A dataset remains unchanged throughout the evaluation process. In contrast, an environment is an interactive simulation that represents a real-world scenario of interest. A GUI environment includes the GUI interface of a mobile phone or a desktop. Unlike datasets, environments are dynamic, actions taken within the environment can alter its state, hence, allowing modeling the problem as Markov Decision Processes (MDPs) or Partially Observable MDPs (POMDPs), with defined action, state, and observation spaces, and a state transition function.
@@
We present a summary of existing GUI agent benchmarks in Table 1.
@@ 3.1 Static Datasets
3.1.1 Closed-World Datasets.
@@
Evaluation on static open-world datasets balances the ease of evaluation with realism since the agents interact with real-world websites. However, due to the nature of real-world websites, they are often unpredictable and prone to changes, which makes it more challenging to reproduce and compare with prior methods.
@@ 3.2 Interactive Environments
Closed-world interactive environments provide controlled and reproducible settings for evaluating agent capabilities. MiniWoB offers synthetic web tasks requiring interactions with webpages using mouse and keyboard inputs (Shi et al., 2017). It focuses on fundamental skills like button clicking and form filling, providing a baseline for evaluating low-level interaction. CompWoB extends MiniWoB with compositional tasks, requiring agents to handle multi-step workflows and generalize across task sequences (Furuta et al., 2023). This introduces dynamic dependencies that reflect real-world complexity. WebShop simulates e-shopping tasks that challenge agents to navigate websites, process instructions, and make strategic decisions (Yao et al., 2022). WebArena advances realism with self-hosted environments across domains like e-commerce and collaborative tools, requiring agents to manage long-horizon tasks (Zhou et al., 2023b). VisualWebArena adds multimodal challenges, integrating visual and textual inputs for tasks like navigation and object recognition (Koh et al., 2024a). Shifting to enterprise settings, WorkArena evaluates agent performance in complex UI environments, focusing on knowledge work tasks in ServiceNow platform (Drouin et al., 2024). WorkArena++ extends this benchmark by introducing more challenging tasks (Boisvert et al., 2024). ST-WebAgentBench incorporates safety and trustworthiness metrics, assessing policy adherence and minimizing risky actions, critical for business deployment (Levy et al., 2024). VideoWebArena introduces long-context video-based tasks, requiring agents to understand instructional videos and integrate them with textual and visual data to complete tasks. It emphasizes memory retention and multimodal reasoning (Jang et al., 2024). In simulated desktop environments, OSWorld (Xie et al., 2024) provides the first realistic operating system setting for evaluating multimodal GUI agents. Spider2-V (Cao et al., 2024) builds on this direction by targeting professional-level data science and engineering tasks. BrowserGym (Chezelles et al., 2024) develops a unified ecosystem consisting of seven web agent benchmarks for developing and evaluating web agents.
@@ 3.3 Evaluation Metrics
3.3.1 Task Completion Metrics.
@@
The majority of benchmarks use task completion rate as the primary metric to measure GUI agents’ performance. However, different papers define task completion differently. Success can be defined as whether an agent successfully stops at a goal state (Chen et al., 2024; Zhou et al., 2023b), with Zhou et al. (2023b) programmatically checking if the intended outcome has been achieved (e.g., a comment has been posted, or a form has been completed), or whether the returned results exactly match the ground truth labels (Shi et al., 2017; Yao et al., 2022; Koh et al., 2024a; Drouin et al., 2024; Levy et al., 2024; Mialon et al., 2023). Another approach is to measure success based on whether an agent completes all required subtasks (Lai et al., 2024; Zhang et al., 2024e; Pan et al., 2024; Furuta et al., 2023; Jang et al., 2024; Cheng et al., 2024b). This approach can be further extended to measure partial success, as shown in Zhang et al. (2024e). WebVoyager uses GPT-4V to automatically determine success based on the agent’s trajectory, reporting a high agreement rate of 85.3% with human judgments (He et al., 2024b). Instead of using a single final-state success metric, WebLINX measures an overall success rate based on aggregated turn-level success metrics across tasks (Lù et al., 2024). These turn-level metrics, including Intersection over Union and F1, are computed based on the type of action taken. Lastly, there are task-specific metrics to measure success, e.g., using ROUGE-L, F1 for open-ended generation (Liu et al., 2024b; Xu et al., 2024; Wornow et al., 2024), accuracy for multiple choice question tasks (Liu et al., 2024b), Precision and Recall for Standard Operating Procedure validation (Wornow et al., 2024).
3.3.2 Intermediate Step Metrics.
@@
While the task completion rate is a single straightforward metric that simplifies evaluation, it fails to provide clear insights into their specific behaviors. Although some fine-grained metrics measure step-wise performance, their scope remains limited. WebCanvas (Pan et al., 2024) evaluates step scores using three distinct targets: URL Matching, which verifies whether the agent navigated to the correct webpage; Element Path Matching, which checks if the agent interacted with the appropriate UI element, such as a button or text box; and Element Value Matching, which ensures the agent inputted or extracted the correct values, such as filling a form or reading text. WebLINX (Lù et al., 2024) uses an intent match metric to assess whether the predicted action’s intent aligns with the reference intent. Similarly, Mind2Web (Deng et al., 2023) and MT-Mind2Web (Deng et al., 2024) evaluate Element Accuracy by measuring the rate at which the agent selects the correct elements. These systems also measure the precision, recall, and F1 score for token-level operations, such as clicking or typing, and calculate the Step Success Rate, which reflects the proportion of individual task steps completed correctly. While step-wise evaluations provide more fine-grained insight into the agent’s performance, it is often challenging to collect reference labels at the step level while also providing enough flexibility to consider different paths to achieve the original tasks.
3.3.3 Efficiency, Generalization, Safety and Robustness Metrics.
@@ 4 GUI Agent Architectures
This section focuses on various architectural designs of a GUI agent, which we categorize into four main types: (1) Perception: designs that enable the GUI agent to perceive and interpret observations from its environment; (2) Reasoning: designs related to the cognitive processes of a GUI agent, such as using an external knowledge base for long-term memory access or a world model of the environment to support other modules like planning; (3) Planning: designs related to decomposing a task into subtasks and creating a plan for their execution; and (4) Acting: mechanisms that allow the GUI agent to interact with the environment, including representing actions in natural language using specific templates, JSON, or programming languages as action representations. We present a taxonomy of GUI agent architectures in Figure 1.
@@ 4.1 Perception
Unlike API-based agents that process structured, program-readable data Xu et al. (2025b), GUI agents must perceive and understand the on-screen environment that is designed for human consumption. This requires carefully chosen interfaces that allow agents to discover the location, identity, and properties of the interactive elements. Broadly, these perception interfaces can be categorized into four types: accessibility-based, HTML/DOM-based, screen-visual-based, and hybrid ones, with each offering different capabilities and posing distinct privacy and implementation considerations.
@@ 4.1.3 Screen-visual-based Interfaces
With advances in computer vision and multimodal LLMs, agents can utilize screen-visual information, such as screenshots, to perceive the on-screen environment. OmniParser (Lu et al., 2024) utilizes an existing multimodal LLM (e.g., GPT-4V) to parse a screenshot into a structured representation of the UI elements. TAG (Xu et al., 2025a) leverages the inherent attention patterns in pretrained MLLMs to improve GUI grounding without the need for additional fine-tuning. Cradle (Tan et al., 2024), instead of relying on a single screenshot, processes a video recording (i.e., a sequence of screenshots) to enable more general-purpose computer control. However, screen-visual-based perception introduces privacy concerns since entire screenshots may contain sensitive information. Additionally, computational overhead increases as models must handle high-dimensional image inputs. Despite these challenges, such interfaces are crucial for agents operating in environments where high-quality accessibility interfaces and DOM information are unavailable, or environments where dynamic or visual information is crucial, like image or video editing software. A key advantage of this approach is that it requires no application instrumentation, enabling direct deployment across a wide range of applications.
@@ 4.1.4 Hybrid Interfaces
To achieve robust and flexible performance across diverse environments, many GUI agents employ a hybrid approach (Gou et al., 2024; Wu et al., 2024b; Kil et al., 2024). These systems combine accessibility APIs, DOM data, and screen-visual information to form a more comprehensive understanding of the interface. Leading methods in GUI agent tasks, such as OS-Atlas (Wu et al., 2024b) and UGround (Gou et al., 2024), demonstrate that hybrid interfaces that combine visual and textual inputs can enhance performance. Such approaches also facilitate error recovery, when accessibility or DOM data are incomplete or misleading, the agent can fall back on screen parsing, and vice versa.
@@ 4.3 Planning
Planning involves decomposing a global task into multiple subtasks that progressively approach the goal state starting from an initial state Huang et al. (2024). Traditional planning methods, such as symbolic approaches (Kautz and Selman, 1992) and reinforcement learning (Sutton and Barto, 1998), have significant limitations: symbolic methods require extensive human expertise to define rigid system rules and lack error tolerance Belta et al. (2007); Pallagani et al. (2022), while reinforcement learning demands impractical volumes of training data, often derived from costly environmental interactions Acharya et al. (2023). Recent advancements in LLM-powered agents offer a transformative alternative by positioning LLM-powered agents as the cognitive core for planning agents Huang et al. (2024). When equipping agents with GUIs as the medium, LLM-powered agents can directly interact with nearly all application domains and resources to enhance planning strategies. Based on what application domains/resources agents use for planning, we divide existing works into planning with internal and external knowledge.
@@ 4.3.1 Planning with Internal Knowledge
Planning with internal knowledge of GUI agents is to leverage the inherent knowledge to reason and think about the potential plans to fulfill the global task goals Schraagen et al. (2000). WebDreamer (Gu et al., 2024) uses LLMs to simulate the outcomes of the actions of each agent and then evaluates the result to determine the optimal plan at each step. MobA (Zhu et al., 2024) devises a two-level architecture to power the mobile phone management, with a high level for understanding user commands, tracking history memories and planning tasks, and a low level to act the planned module. Agent S (Agashe et al., 2024) introduces an experience-augmented hierarchical planning to perform complex computer tasks.